How strong is the electric field between the plates of a 0.60-?Fair-gap capacitor if they are 2.0mm apart and each has a charge of 71?C ?

Express your answer using two significant figures and include the appropriate units.
Answer #1
C = Q / V
V = Q / C = 72?C / 0.60?F = 120V

E = V / d = 120V / 2.0mm = 60kV/m

Source:

http://physicsguide.blogspot.ca/2006/02/electricity-and-magnetism-tutorial-2.html
Answer #2

C = Q / V
V = Q / C = 71?C / 0.60?F =118V

E = V / d = 118V /2.0mm = 59kV/m